2. (10 points) Consider the following computational problems:

EQDF A = {hA, Bi | A and B are DFAs and L(A) = L(B)}

SUBDF A = {hA, Bi | A and B are DFAs and L(A) ⊆ L(B)}

DISJDF A = {hA, Bi | A and B are DFAs and L(A) ∩ L(B) = ∅}.

Prove that SUBDF A and DISJDF A are each Turing-decidable.

You may (and should) use high-level descriptions of any Turing machines you define. Make sure to provide both a machine definition and a proof of correctness.

2. (10 points) Consider the following computational problems: EQDFA = {(A,B) | A and B are DFAs and L(A) = L(B)) SUBDFA = {(A,B) | A and B are DFAs and L(A) L(B)) DISJDFA = {(A,B) | A and B are DFAs and L(A) L(B) = } Prove that SUBDFA and DISJDFA are each Turing-decidable. You may (and should) use high-level descriptions of any Turing machines you define. Make sure to provide both a machine definition and a proof of correctness.
Answer #1

Solution for SUB_{DFA} is Turing decidable or not.

Given \item[] $SUB_{DFA} = \{ \langle A, B \rangle \st \text{$A$ and $B$ are DFAs and $L(A) \subseteq L(B)$} \}$

Let X and Y be two sets. Then X \subseteq Y iff X \cap \overline{Y} = \Phi

Thus based on the above Idea:

Define TM M_{1} by : M_{1} = "On Input \left \langle A, B \right \rangle where A, BDFA's

  1. Construct a new DFAD from A, B such that DFA   D is intersection of A and \overline{B} as we know that intersection and complement of DFA is decidable. Thus
    L(D) = L(A) \cap \overline{L(B)}
  2. We know that E_{DFA} is decidable. ie Empty DFA is Turing decidable.
    Thus TM M_{1} accepts if L(D) is empty and rejects otherwise.

Solution (b) To prove DISJ_{DFA} is Turing decidable.
$DISJ_{DFA} = \{ \langle A, B \rangle \st \text{$A$ and $B$ are DFAs and $L(A) \cap L(B) = \emptyset$} \} $.

From Demorgan's law we know the fact that if X and Y are two sets

Then X\capY = \overline{\overline{X}\cup \overline{Y}} . And we also know that union and complement of DFA are decidable thus the intersection is also decidable.

Based on Above Idea:

Define TM M_{2} by : M_{2} = "On Input \left \langle A, B \right \rangle where A, BDFA's

  1. Construct a new DFAD from A, B such that DFA   D is intersection of A and B as we know that intersection of two DFAs is decidable. Thus
    L(D) = L(A) \cap L(B)
  2. We know that E_{DFA} is decidable. ie Empty DFA is Turing decidable.
    Thus TM M_{2} accepts if L(D) is empty and rejects otherwise.

Proof of correctness is just to tell that M1 and M2 is decider and the L(M1) and L(M2) are indeed equivalent to SUB_DFA and DISJ_DFA.